Quantitative Problem: Bellinger Industries is considering two projects for inclusion in its capital budget, and you have been asked to do the analysis. Both projects' after-tax cash flows are shown on the time line below. Depreciation, salvage values, net operating working capital requirements, and tax effects are all included in these cash flows. Both projects have 4-year lives, and they have risk characteristics similar to the firm's average project. Bellinger's WACC is 10%.

0 1 2 3 4
Project A -1,100 700 420 250 300
Project B -1,100 300 355 400 750

What is Project A's payback? Do not round intermediate calculations. Round your answer to four decimal places.

__________years

What is Project A's discounted payback? Do not round intermediate calculations. Round your answer to four decimal places.

___________years

What is Project B's payback? Do not round intermediate calculations. Round your answer to four decimal places.

__________years

What is Project B's discounted payback? Do not round intermediate calculations. Round your answer to four decimal places.

_____________years
Answer #1

D15 fix =2+116.53/187.83 AB C D 1 A 2 year Cash flows pv@10% Present value Cumulative Cash flows Discounted Cumulative Cash F

D15 fax =2+116.53/187.83 А в E pv@10% 1A 2 year 30 4 1 5 2 63 Cash flows -1100 700 420 250 300 =IRR(B3:37) =C3/1.1 =C4/1.1 C5

LIPDUdlu PUPIL Algieril Number 015 - fx =3+233.36/512.26 B 1 B 2 year Cash flows pv@10% Present value Cumulative Cash flows D

D15 fc =3+233.36/512.26 pv@10% Cash flows -1100 300 1 B 2 year 30 4 1 52 63 7 4 8 IRR 9 Present value =B3*C3 =B4*04 =B5*C5 =B